OpenStudy (anonymous):

evaluate the given integral using a substitution 7.4e^(-4x+7)dx

OpenStudy (anonymous):

\[ u(x) = -4x+7\\ \frac{du}{dx} = -4 \\ du = -4dx\\ dx= -\frac{du}{4} \] So then \[\Large 7.4e^{-4x+7}dx\to 7.4e^{u}\left(-\frac{du}{4}\right) \]

OpenStudy (anonymous):

Ok that is a lot easier than I thought. Thank you